# It Is Giving Major Satisfaction: Why Fairness Matters for Software Practitioners

### v2

@@ 1. Introduction
Building on this work, we are interested in exploring perception of fairness as another potential factor that could shape job satisfaction. Similar to Storey et al.’s (Storey et al., 2021) approach in identifying key factors that influence job satisfaction, our study seeks to investigate whether and how perceptions of fairness relate to job satisfaction among software practitioners, both as a collective, and from different backgrounds.
2. Research Questions
To address whether and how perceptions of fairness relate to job satisfaction, we conducted an online survey, receiving a total of 108 responses. We used the 20-item fairness scale developed by Colquitt (Colquitt, 2001) to measure fairness perceptions across four dimensions: distributive, procedural, interpersonal, and informational fairness. Job satisfaction was measured using 15 satisfaction factors inspired by Storey et al. (Storey et al., 2021).
The survey results indicate a mixed perception of fairness across different dimensions among software practitioners. While interpersonal fairness is generally perceived positively, there seems to be moderate satisfaction with distributive, procedural and informational fairness. In terms of job satisfaction, most respondents expressed satisfaction, though there was notable dissatisfaction with promotion opportunities.
After observing software practitioners’ perceptions of fairness and job satisfaction, our next step is to determine whether and how these perceptions of fairness shape practitioners’ job satisfaction.
RQ1: Does fairness perception shape job satisfaction among software practitioners? If so, how?
Our regression analysis shows that perceptions of fairness significantly shape job satisfaction among software practitioners. Fairness across all four dimensions; distributive, procedural, interpersonal, and informational, significantly enhances overall job satisfaction and job security for software practitioners.
While these insights provide a general understanding of how fairness perceptions shape job satisfaction, our survey results also highlight notable differences based on demographic factors. Respondents with 3–5 years of experience generally reported more positive perceptions of fairness than those with over 7 years of experience. Similarly, respondents from diverse teams expressed more favorable fairness perceptions than those whose ethnicity is underrepresented in their team. These differences indicate that perceptions of fairness and their impact on job satisfaction may vary significantly across different groups.
This observation aligns with existing literature that highlights demographic disparities in the SE industry, which is predominantly young, White, and male. Studies have shown that minority groups often face unique challenges. For example, Trinkenreich et al. identified that women in the software industry encounter issues like work-life balance, the glass ceiling, and sexism (Trinkenreich et al., 2022a). Canedo et al. found that women are underrepresented in software development and open source projects, often due to male-dominated workplace conditions (Canedo et al., 2019). Furthermore, Nadri et al. reported that less than 10% of contributions to open source projects come from visibly Non-White developers (Nadri et al., 2021). Similarly, Baltes et al. discuss age-related biases, noting that older developers are often perceived as less adaptable, which can limit their career progression (Baltes et al., 2020).
Given these differences, it is possible that the majority’s perspective does not necessarily reflect the experiences of the minority. By examining these perceptions across various demographics, we can ensure that we capture a diverse range of experiences. This is important for understanding what aspects of fairness shape job satisfaction across the SE community, ensuring that the insights developed are inclusive and address the needs of all practitioners.
RQ2: Does the relationship between fairness perceptions and job satisfaction change based on demographic groups? If so, how?
Through moderation analysis, we found that the relationship between fairness perceptions and job satisfaction does indeed vary across demographic groups. This relationship is found to be stronger for female, underrepresented, less experienced practitioners, and those with work limitations.
While RQ1 and RQ2 reveal that perceptions of fairness are linked to job satisfaction and that these relationships differ across demographic groups, it remains unclear how to apply these insights practically in the workplace. Understanding that, for example, interpersonal fairness influences overall job satisfaction is important, but it does not specify what actions can be taken to address or improve these perceptions in specific scenarios.
Building on the work of Sesari et al. (Sesari et al., 2024), who identified various scenarios in which fairness problems persist within SE contexts, we sought to explore these scenarios further. In our survey, respondents were asked to rate the extent to which they have experienced fairness across 10 SE contexts, adapted from the study by Sesari et al. (Sesari et al., 2024). Our analysis found that most respondents experienced unfairness in the context of demand (e.g., pressing deadline), authorship (e.g., code attribution), income (e.g., being underpaid), and allocation of work (e.g., assigned tasks unsuited to skills).
Given these findings, it is important to examine how these fairness issues manifest in specific SE contexts and determine which scenarios most significantly affect job satisfaction. This approach can help translate the broader understanding of fairness perceptions into actionable strategies that address specific SE situations.
RQ3: How do fairness perceptions influence job satisfaction in specific SE contexts?
Fairness experiences across all four dimensions in contexts such as authorship significantly enhance overall job satisfaction and job security among software practitioners. Practices related to recognizing contributions, giving credit, and ensuring that the right individuals are acknowledged for their work offer opportunities to implement fairness.
Building on our findings from RQ2, where we observed that the relationship between fairness perceptions and job satisfaction varies across demographic groups, it becomes essential to explore this dynamic in specific SE scenarios. Given the insights from RQ3, it is also important to examine how these manifestations of fairness issues differ across demographic groups within the SE context.
For instance, if we find that being treated politely, with respect, and dignity during evaluations significantly affects job satisfaction, we should investigate which demographic groups this applies to the most. Is it more impactful for female developers, suggesting that gender plays a key role in this relationship?
RQ3.1: How does fairness perceptions influencing job satisfaction differ across demographic groups in the SE context?
Our analysis confirms that fairness perceptions influencing job satisfaction do indeed differ across demographic groups within SE contexts. Female developers are more impacted by fairness in policy implementation and evaluations. Underrepresented practitioners value fairness in policy implementation and during high-demand situations. Less experienced practitioners place greater emphasis on fairness when navigating working hours, while practitioners with work limitations prioritize fairness during demanding situations.
@@ 3.3. How Fairness and Job Satisfaction are Related
4. Survey Design
The present study seeks to investigate how perceptions of fairness relate to job satisfaction among software practitioners. To achieve this, we focus on two key constructs: perceptions of fairness and job satisfaction. The measurement of these constructs has been comprehensively and exhaustively developed and tested using survey methods, making a survey the appropriate tool for our study.
Additionally, to explore the specific real-world scenarios in which these fairness aspects are most likely to affect job satisfaction (RQ3), we ask participants how they have experienced fairness-related contexts identified in the SE field.
@@ 4. Survey Design
4.1. Measuring Perception of Fairness
@@ 4.2. Measuring Job Satisfaction
Consequently, we refined our scope from the original 44 factors to 15 that are more closely aligned with fairness, supported by empirical evidence. The selected factors, presented in Table 2, allow us to analyze job satisfaction in a way that is directly linked to perceptions of fairness.
@@
4.3. Measuring Fairness Experiences in SE Contexts
@@
A study by Sesari et al. (Sesari et al., 2024) identifies ten distinct contexts where fairness is a concern within the SE field: income, treatment, allocation of work, recruitment, working hours, demand, evaluation, authorship, policy, and customer/client relations. These contexts represent primary areas where fairness challenges are most likely to arise, offering valuable insights into where software engineers perceive fairness issues in their daily professional experiences. To investigate these contexts further, we asked survey respondents to rate the extent to which they have experienced fairness in each of the ten contexts identified by Sesari et al. (Sesari et al., 2024). This approach allows us to systematically measure the extent to which fairness perceptions are experienced in various SE scenarios.
4.4. Collecting Demographic Information
@@
Gender: Male, Female. Given the historical gender biases in the SE field, particularly in areas like hiring and promotion (Wang and Redmiles, 2019), we anticipate that gender may play a significant role in shaping job satisfaction.
@@ 4.5. Respondent Recruitment
5. Research Methods
In this section, we detail the methodologies employed to answer the research questions posed in this study. The scripts for all steps are provided in the supplementary material (fig, 2024).
5.1. Pre-processing
@@
The shortest response time was 0.6 minutes, which we deemed too short to yield meaningful data given the expected 10-minute duration. Research supports that very quick survey completions can indicate less attentive responses (Wood et al., 2017). On reviewing the 0.6-minute response, we found missing answers in key areas, reinforcing concerns about data quality for very fast responses. We also examined responses with durations beyond the third quartile (Q3), ranging from 17 minutes to $\sim$6 days. The extended response times might be due to allowing respondents to return to the survey later, as their progress remains saved even if the tab is left open. Therefore, we did not find it necessary to exclude these, as longer completion times do not necessarily correlate with lower data quality.
@@
To ensure the reliability of our survey, we also conducted an Omega reliability test (McDonald, 1970), which is well-suited for multidimensional surveys. Omega reliability test is based on a confirmatory factor analysis model that recognizes the possibility of multiple factors or dimensions within a survey (McDonald, 1970). In our survey, there are items representing job satisfaction, items representing fairness, and items representing fairness contexts. Given the multidimensional nature of these items, Omega reliability test was the appropriate choice to capture the varying contributions of these items across different dimensions.
The Omega reliability test conducted before and after imputation showed a reliability score of 0.95, indicating that 95% of the variance in our survey responses is attributable to consistent, reliable information. The consistency of these values before and after imputation confirms that our method for handling missing data did not compromise the survey’s overall reliability. After pre-processing, we retained 108 responses for further analysis.
@@ 5.3. Grouping Job Satisfaction Factors
5.4. Ordinal Logistic Regression Analysis
After establishing the job satisfaction clusters, we aimed to understand how fairness perception or fairness contexts are associated with the probability of job satisfaction being in a higher or lower category. To achieve this, we performed an ordinal logistic regression analysis. Ordinal logistic regression predicts the probability of an observation falling into one of several ordered categories, making it ideal for analyzing outcomes like job satisfaction levels (Agresti, 2010).

### v3

@@ 1. Introduction
Building on this work, our study aims to explore the perception of fairness as a critical factor influencing developer job satisfaction. Based on prior organizational behavior research, we theorize that fairness perceptions influence job satisfaction. Colquitt’s meta-analytic review (Colquitt et al., 2001) found that perceptions of distributive, procedural, interpersonal, and informational fairness each significantly predict job satisfaction across industries. This relationship has been supported in diverse settings (Judge and Colquitt, 2004; Konovsky and Cropanzano, 1991; Moorman, 1991). Drawing on these findings, and consistent with emerging SE research highlighting the importance of fairness issues (German et al., 2018; Sesari et al., 2024), we posit that developers who perceive higher fairness across these dimensions will report higher job satisfaction. Thus, fairness perceptions are conceptualized as key antecedents of job satisfaction in the SE context.
While previous studies in SE have separately explored fairness-related challenges (Sesari et al., 2024; German et al., 2018) and job satisfaction factors (Fagerholm et al., 2014; Storey et al., 2021), no study has directly linked developers’ fairness perceptions to their job satisfaction. We address this critical gap by empirically investigating how fairness perceptions shape developers’ satisfaction with their work. Through a survey among software practitioners, we measure fairness perceptions based on Colquitt’s organizational justice dimensions and analyze their relationship with job satisfaction outcomes. Our approach is consistent with recent SE research that emphasizes grounding empirical investigations in clear theoretical constructs and modeling perception-to-outcome pathways (e.g., (Verwijs and Russo, 2024; Lambiase et al., 2024)). Our findings aim to inform future empirical studies or organizational assessments focused on fairness-related workplace outcomes in SE and provide practical implications for fostering fair, satisfying work environments that support talent retention and developer well-being.
The rest of the paper is structured as follows: Section 2 provides background on fairness and job satisfaction in SE. Section 3 presents the conceptual foundations by defining the key constructs and explaining the hypothesized relationships behind our research questions. Section 4 outlines the methodology, including data collection, data processing and analysis strategies. Section 5 presents the empirical findings. Section 6 discusses the implications of these findings. Section 7 outlines threats to validity, and Section 8 concludes the paper with key takeaways and future research directions.
@@ 2.3. How Fairness and Job Satisfaction are Related
3. Conceptual Framework
In this section, we present the conceptual foundations of our study by defining the constructs and explaining the hypothesized relationships between them. We ground our framework in established organizational fairness and job satisfaction theories, while extending their application to SE.
3.1. Theoretical Constructs
Our study centers on three major constructs: fairness perceptions, job satisfaction, and fairness experiences in SE contexts.
Fairness Perceptions. We adopt Colquitt’s multidimensional framework of organizational fairness (Colquitt, 2001), which defines fairness across four dimensions: distributive, procedural, interpersonal, and informational fairness.
Job Satisfaction. To assess job satisfaction in the SE context, we draw on the framework by Storey et al. (Storey et al., 2021), which outlines 44 job satisfaction factors relevant to software developers.
Fairness Experiences in SE. We also explore fairness perceptions across specific SE scenarios. Drawing on prior work by Sesari et al. (Sesari et al., 2024), we included items that capture the extent of fairness challenges in SE, such as unfair allocation of work, unfair income distribution, and unfair recruitment.
3.2. Framing the Research Questions
Based on prior organizational behavior research, we theorize that fairness perceptions influence job satisfaction. Colquitt’s meta-analytic review (Colquitt et al., 2001) found that perceptions of distributive, procedural, interpersonal, and informational fairness each significantly predict job satisfaction across industries (Judge and Colquitt, 2004; Konovsky and Cropanzano, 1991; Moorman, 1991). Drawing on these findings and emerging SE research on fairness (German et al., 2018; Sesari et al., 2024), we posit that developers who perceive higher fairness across these dimensions will report higher job satisfaction. Thus, we formulate the following research questions:
RQ1::
Does fairness perception shape job satisfaction among software practitioners? If so, how?
While RQ1 explores the general relationship between fairness perceptions and job satisfaction, existing literature also highlights demographic disparities in SE, a field predominantly young, White, and male. Studies have shown that minority groups often face unique challenges. For example, Trinkenreich et al. (Trinkenreich et al., 2022a) identified that women in the software industry encounter issues such as work-life balance, the glass ceiling, and sexism. Canedo et al. (Canedo et al., 2019) found that women are underrepresented in software development and open source projects, often due to male-dominated workplace conditions. Nadri et al. (Nadri et al., 2021) reported that less than 10% of contributions to open source projects come from visibly non-White developers. Similarly, Baltes et al. (Baltes et al., 2020) discussed age-related biases, noting that older developers are often perceived as less adaptable, which can hinder career progression.
Given these disparities, we theorize that the majority’s perspective may not fully capture the experiences of minority groups. By examining perceptions across diverse demographics, we can ensure that our findings reflect a broader range of experiences.
RQ2::
Does the relationship between fairness perceptions and job satisfaction change based on demographic groups? If so, how?
While RQ1 and RQ2 reveal important patterns, they do not explain how fairness perceptions operate in specific SE contexts. Building on the work of Sesari et al. (Sesari et al., 2024), who identified key scenarios in which fairness problems persist in SE, we further investigate how such issues manifest and impact job satisfaction in situational settings.
RQ3::
How do fairness perceptions influence job satisfaction in specific SE contexts?
Building on RQ2 and RQ3, we also explore how the relationship between fairness perceptions and job satisfaction in SE contexts varies across demographic groups. For instance, if respectful treatment during evaluations affects job satisfaction, which demographic groups are most impacted? Is this effect stronger for female developers, suggesting gender plays a critical role?
RQ3.1::
How does fairness perception influencing job satisfaction differ across demographic groups in SE contexts?
4. Methodology
The present study seeks to investigate how perceptions of fairness relate to job satisfaction among software practitioners (RQ1). To achieve this, we focus on two key constructs: perceptions of fairness and job satisfaction. The measurement of these constructs has been comprehensively and exhaustively developed and tested using survey methods, making a survey the appropriate tool for our study. Additionally, to explore the specific real-world scenarios in which these fairness aspects are most likely to affect job satisfaction (RQ3), we ask participants how they have experienced fairness-related contexts identified in the SE field.
@@ 4. Methodology
Finally, we detail the steps employed to answer the research questions posed in this study. The scripts for all steps are provided in the supplementary material (fig, 2024).
4.1. Data Collection
4.1.1. Measuring Perception of Fairness
@@ 4.1.2. Measuring Job Satisfaction
Consequently, we refined our scope from the original 44 factors to 15 that are more closely aligned with fairness, supported by empirical evidence. The complete list of 44 factors is available in the original study by Storey et al. (Storey et al., 2021), and we provide an overview of excluded factors in the supplementary material (fig, 2024). The selected factors, presented in Table 2, allow us to analyze job satisfaction in a way that is directly linked to perceptions of fairness.
@@
4.1.3. Measuring Fairness Experiences in SE Contexts
@@
A study by Sesari et al. (Sesari et al., 2024) identifies ten distinct contexts where fairness is a concern within the SE field: income, treatment, allocation of work, recruitment, working hours, demand, evaluation, authorship, policy, and customer/client relations. These contexts represent primary areas where fairness challenges are most likely to arise, offering valuable insights into where software engineers perceive fairness issues in their daily professional experiences. To investigate these contexts further, we asked survey respondents to rate the extent to which they have experienced fairness in each of the ten contexts identified by Sesari et al. (Sesari et al., 2024). This approach allows us to systematically measure the extent to which fairness perceptions are experienced in various SE scenarios. Below is the list of fairness experiences in SE contexts faced by practitioners, adapted from the study by Sesari et al. (Sesari et al., 2024):
  • income: fairness in how practitioners are compensated for their contributions;
  • treatment: fairness in how practitioners are treated and how they treat others in the workplace;
  • allocation of Work: fairness in how tasks and responsibilities are distributed among team members;
  • recruitment: fairness in the processes used to hire or onboard practitioners;
  • working hours: fairness in expectations around the amount and distribution of working time;
  • demand: fairness in the expectations or requests placed on practitioners by those in authority;
  • evaluation: fairness in how work performance or contributions are assessed;
  • authorship: fairness in how credit is assigned for contributions to code, design, or other outputs;
  • policy: fairness in organizational policies or rules that affect practitioners’ work;
  • customer/client relations: fairness in interactions between practitioners (or their organizations) and external stakeholders such as clients or customers.
4.1.4. Collecting Demographic Information
@@
Gender: Male, Female. Given the historical gender biases in the SE field, particularly in areas like hiring and promotion (Wang and Redmiles, 2019), we anticipate that gender may play a significant role in shaping job satisfaction. While we acknowledge that broader or open-text gender categories could capture more diverse identities, such an approach was beyond the scope of this study and would have posed challenges for statistical subgroup analysis given the sample size. Future research could specifically investigate fairness experiences of gender-diverse individuals in SE through targeted sampling and mixed-methods designs.
@@ 4.1.5. Respondent Recruitment
We applied purposive sampling (Baltes and Ralph, 2022), recruiting respondents who are professional software practitioners via LinkedIn, social media, and professional events. This approach enabled us to reach individuals likely to provide relevant insight into fairness and job satisfaction in software development contexts. Following Baltes and Ralph’s classification (Baltes and Ralph, 2022), our sampling strategy corresponds to a form of convenience sampling with elements of expert sampling, as we selectively targeted working professionals in software engineering through professional networks and domain-relevant channels.
@@
4.1.6. Ethical Considerations
This study was approved by the relevant ethics committees at the authors’ institutions under approval numbers 97126442 and UCL/CSREC/R/40. All participants were informed of the purpose of the study, their rights, and the handling of their data through a participant information sheet and informed consent form, both of which were presented before the survey began. Participants provided their explicit consent before proceeding with the survey. Details about data storage, anonymity, and the handling of sensitive information were provided in the Participant Information Sheet, which is included in the supplementary materials (fig, 2024). The informed consent form is also available.
4.2. Pre-processing and Assessing Survey Data Quality
@@
We implemented response time analysis as a proxy for detecting inattentive or careless responses, consistent with recommendations in survey methodology literature (Read et al., 2022). Responses completed in less than one minute were flagged as likely inattentive. While explicit attention-check questions are common in some survey designs, we chose not to include them to maintain engagement and trust among respondents. Prior research indicates that experienced participants may recognize such checks, potentially reducing their effectiveness and introducing bias (Pei et al., 2020; Almog et al., 2023).
We established a threshold to detect outliers, focusing on extremely short response times, such as responses completed in less than one minute. Using the Interquartile Range (IQR) method, we identified responses with durations in the first quartile as potential outliers. The shortest response time was 0.6 minutes, which we deemed too short to yield meaningful data given the expected 10-minute duration. Research supports that very quick survey completions can indicate less attentive responses (Wood et al., 2017). On reviewing the 0.6-minute response, we found missing answers in key areas, reinforcing concerns about data quality for very fast responses. We also examined responses with durations beyond the third quartile (Q3), ranging from 17 minutes to $\sim$6 days. The extended response times might be due to allowing respondents to return to the survey later, as their progress remains saved even if the tab is left open. Therefore, we did not find it necessary to exclude these, as longer completion times do not necessarily correlate with lower data quality.
@@
To ensure the validity and reliability of our survey constructs, we conducted both Exploratory Factor Analysis (EFA) and Confirmatory Factor Analysis (CFA), followed by an assessment of internal consistency using the Omega reliability coefficient (McDonald, 1970).
The EFA results aligned well with theoretical expectations. For fairness perceptions (20 items), a four-factor structure emerged, corresponding to Colquitt’s dimensions of distributive, procedural, interpersonal, and informational fairness. The model demonstrated strong fit (RMSEA = 0.059, TLI = 0.937, RMSR = 0.04), with items loading cleanly onto their respective factors and minimal cross-loadings. For job satisfaction (16 items), EFA revealed acceptable fit (RMSEA = 0.082, TLI = 0.873, RMSR = 0.05), consistent with conceptual groupings adapted from Storey et al. (Storey et al., 2021). For fairness contexts (10 items), parallel analysis suggested a unidimensional structure, which was supported by good fit indices (RMSEA = 0.081, TLI = 0.903, RMSR = 0.07), indicating that fairness experiences across SE scenarios reflect a single construct.
CFA was subsequently performed to confirm these structures. The fairness perceptions model showed good fit (CFI = 0.921, TLI = 0.909, RMSEA = 0.076, SRMR = 0.076), supporting Colquitt’s four-dimensional framework. The job satisfaction model demonstrated acceptable fit (CFI = 0.857, TLI = 0.830, RMSEA = 0.101, SRMR = 0.082), reflecting the exploratory nature of adapting Storey et al.’s factors. The fairness contexts model confirmed the unidimensional structure with strong fit indices (CFI = 0.919, TLI = 0.896, RMSEA = 0.087, SRMR = 0.061).
Across all constructs, item-level standardized factor loadings were generally high, ranging from 0.52 to 0.97 for fairness perceptions, 0.49 to 0.87 for job satisfaction, and 0.41 to 0.78 for fairness contexts. These factor loadings indicate that individual items contributed meaningfully to their respective latent constructs.
Finally, to assess internal consistency, we applied the Omega reliability coefficient, which is particularly suited for multidimensional constructs and provides a more accurate estimate of reliability than Cronbach’s alpha when factor loadings vary (McDonald, 1970). The Omega values were strong across all constructs, with 0.94 for fairness perceptions, 0.93 for job satisfaction, and 0.89 for fairness contexts. These results indicate a high level of reliability, which confirm that the variance in responses is largely attributable to consistent underlying constructs. Notably, Omega values remained stable before and after multiple imputation, which demonstrate that our handling of missing data did not compromise reliability. After pre-processing, we retained 108 responses for analysis.
As mentioned in Section 4.1.5, we employed purposive sampling focused on recruiting software practitioners from diverse backgrounds to align our sampling strategy with our study’s exploratory goals. While our final sample includes 108 respondents, this size is appropriate for our research objective: examining theorized relationships between fairness perceptions and job satisfaction, rather than producing population-level estimates. As Baltes and Ralph emphasize in their critical review of sampling in SE research, “there is no universally correct sample size,” and what matters is how well the sampling strategy supports the intended claims of a study (Baltes and Ralph, 2022). They further argue that when a study aims to explore social or organizational constructs, diversity of experience and transparent reporting are often more meaningful than statistical representativeness. Following these recommendations, we prioritized recruiting participants across a range of demographics and clearly report our sampling limitations to avoid overgeneralization.
@@ 4.4. Grouping Job Satisfaction Factors
4.5. Statistical Analysis
4.5.1. Ordinal Logistic Regression Analysis
After establishing the job satisfaction clusters, we aimed to understand how fairness perception or fairness contexts are associated with the probability of job satisfaction being in a higher or lower category. To achieve this, we performed an ordinal logistic regression analysis. We performed an ordinal logistic regression analysis to explore the relationships between survey items, specifically addressing RQ1 and RQ3. Ordinal logistic regression predicts the probability of an observation falling into one of several ordered categories, making it ideal for analyzing outcomes like job satisfaction levels (Agresti, 2010).
@@ 4.5.1. Ordinal Logistic Regression Analysis
To address RQ1, we applied this analysis to examine the relationship between perceptions of fairness and the seven job satisfaction clusters. For instance, if a practitioner perceives high levels of distributive fairness, the ordinal logistic regression might predict a higher likelihood of reporting increased satisfaction in clusters related to compensation (COM). This suggests that perceptions of fair distribution of resources and rewards are linked to how satisfied practitioners feel with their compensation. For RQ3, we aimed to translate the broader understanding of fairness perceptions into actionable strategies by examining how specific fairness perceptions relate to job satisfaction within various SE contexts. To do this, we applied a similar analysis to the one used for RQ1, focusing on the relationship between fairness experiences and job satisfaction in specific scenarios.